OpenStudy (anonymous):

Solve [3/(x-1)] - [(2x+10)/(x^2+2x-3)] = 1/3 help please , thanks!

OpenStudy (anonymous):

\[ [3/(x-1)] - [(2x+10)/(x^2+2x-3)] = 1/3\]

OpenStudy (anonymous):

1/x+3=1/3 x=0

OpenStudy (anonymous):

(3/x-1)*(x+3)....(3x+9-2x-10)/(x-1)(x+3)=1/x+3=1/3...x=0

OpenStudy (anonymous):

\[(3(x+3)-(2x+10))/{(x-1)(x+3)}=1/3\] \[(x-1)/{(x-1)(x+3)}=1/3\] 1/(x+3)=1/3 x=0

OpenStudy (anonymous):

thank you so much, i understand now :)